# EdgeGaussians - 3D Edge Mapping via Gaussian Splatting

### v1

@@ Abstract
With their meaningful geometry and their omnipresence in the 3D world, edges are extremely useful primitives in computer vision. 3D edges comprise of lines and curves, and methods to reconstruct them use either multi-view images or point clouds as input. State-of-the-art image-based methods first learn a 3D edge point cloud then fit 3D edges to it. The edge point cloud is obtained by learning a 3D neural implicit edge field from which the 3D edge points are sampled on a specific level set (0 or 1). However, such methods present two important drawbacks: i) it is not realistic to sample points on exact level sets due to float imprecision and training inaccuracies. Instead, they are sampled within a range of levels so the points do not lie accurately on the 3D edges and require further processing. ii) Such implicit representations are computationally expensive and require long training times. In this paper, we address these two limitations and propose a 3D edge mapping that is simpler, more efficient, and preserves accuracy. Our method learns explicitly the 3D edge points and their edge direction hence bypassing the need for point sampling. It casts a 3D edge point as the center of a 3D Gaussian and the edge direction as the principal axis of the Gaussian. Such a representation has the advantage of being not only geometrically meaningful but also compatible with the efficient training optimization defined in Gaussian Splatting. Results show that the proposed method produces edges as accurate and complete as the state-of-the-art while being an order of magnitude faster. The code is released at https://github.com/kunalchelani/EdgeGaussians.
@@ 1 Introduction
3D edges can be in the form of straight 3D lines or 3D curves, which we will refer to respectively as lines and curves for simplicity. Seminal works reconstruct 3D lines from images with a Structure-from-Motion (SfM) approach where line segments are detected from images, matched, and triangulated into 3D lines [82, 83, 5, 56, 6, 63]. The approach also extends to 3D curves [36, 68, 65, 37]. However, the lack of repeatability in the 2D edge detection and the limited robustness in the edge matching remain performance bottlenecks for such approaches. Recent works, although, have made impressive progress for mapping lines [60, 59, 46].
Figure 1: EdgeGaussians: the proposed method learns 3D edge points with an explicit representation where an edge point is cast as a 3D Gaussian centered at that point and the edge direction defines the main orientation of the Gaussian. Left to right: 2D edge maps generated by off-the-shelf detectors [73, 62] are used as supervisory signals to train the 3D edge Gaussians. Once trained, the Gaussian are clustered based on the spatial proximity of their centers and their direction consistency. Parametric edges are fitted on top of these clusters.
@@
3D edge reconstruction from 3D point clouds is free of those limitations and usually follows three steps: classify which 3D points lie on edges, cluster the points belonging to the same edge, and link them [80, 8, 92, 78]. However, the classification is usually hindered by the extreme imbalance between the edge point and the non-edge points, and noisy point clouds can lead to spurious classification. An efficient alternative is to operate directly on a 3D edge point cloud, i.e., a point cloud with points only on the 3D edges.
@@
Such an edge point cloud can be learned from images or 2D edge maps [18, 73] using neural fields [87, 44, 91] trained to represent the 3D edges. Once trained, 3D edge points are sampled from level sets of the neural fields on top of which the 3D edges are fitted. However, such neural fields are computationally expensive and require long training times: recent methods can take 1h30 [91], up to 14h [87] to train on a simple object from the ABC-NEF [91, 41] dataset of CAD models (see Sec. 4 for precise runtimes). Another limitation is the accuracy at which neural implicit fields can represent 3D edges: in theory, the 3D edge points lie on a level set of the field (0 for surface representations [44, 87] and 1 for volumetric ones [91]). In practice though, sampling points on such exact level sets is not feasible due to limited float precision or noise in the training. To compensate for this, the point sampling is done within an $\epsilon$-bound of the level sets [91, 44] but such points do not lie accurately on the 3D edges, requiring post-processing to correct these errors.
@@
To make the 3D edge reconstruction simpler and more efficient while preserving accuracy, we propose to learn an explicit representation of the 3D edge points. It directly learns an edge point cloud, hence bypassing the need for level-set point sampling and post-processing. It also directly learns the point’s edge direction instead of inferring it from the 3D edge points [44], which makes the 3D edge fitting simpler. Last, it is on par with existing learning-based methods and runs much faster, e.g., 30 times faster than EMAP [44] and 17 times faster than NEF [91]. The proposed method defines a 3D edge point as the center of a 3D Gaussian and the edge direction defines the main direction of the 3D Gaussian. Such a representation is geometrically meaningful and is trained in a fast and straightforward manner by adopting the training optimization defined in Gaussian Splats [38]. The optimization is adapted to accommodate the specificities of 3D edge learning mainly the sparsity and the occlusion of the 3D edges. The training remains simple and is supervised with off-the-shelf 2D edge maps [73, 62], as in [44, 91].
@@ 2 Related work
3D edge reconstruction methods vary based on the data they are generated from, e.g., images, point clouds, and whether they reconstruct only lines, only curves, or both.
@@
The earliest methods reconstruct 3D lines from images with an SfM approach that detects lines [77, 59, 32, 86, 2], matches them across images based on line descriptors [60, 42, 76, 7], and lifts them to 3D with triangulation [83, 82, 95, 5, 56, 6, 13, 94, 53, 46, 63, 66] or epipolar geometry [27, 28, 29]. The main challenges are matching lines robustly and repetitively detecting lines even when their inconsistent endpoints across views due to occlusions. Existing solutions involve leveraging geometric cues in the scene such as planes [81] and Manhattan frames [67, 16] or replacing hand-crafted steps [77, 7, 76] of the pipeline by their learning counterparts [59, 60, 42, 61]. While recent works [46, 27, 28, 29, 26] have made impressive progress in addressing these limitations, the robustness of the line detection and matching remains a performance bottleneck. These limitations also hold for similar methods reconstructing curves-SfM [36, 68, 65, 37]. This motivates alternative approaches to work around line matching by estimating 3D lines from images using geometric graph optimization on the detected line segments [34, 64], or directly predicting the 3D wireframes in a learning end-to-end manner [97, 49]. A 3D wireframe is a set of 3D lines that define the layout of the scene and form a graph with vertices at edge junctions. Most methods address the 3D line or wireframe mapping problem, with fewer contributions to 3D curve mapping. In this work, we address the 3D mapping of both lines and curves.
Another way to work around the problem of reliable line detection and matching is to directly operate on 3D structures such as depth maps [52], meshes [15], and point clouds obtained from laser scanners [48] or SfM [70]. Given a set of 3D points, most approaches first classify which 3D points lie on edges, cluster the points that lie on the same edge, link them, and fit a parametric edge to each cluster. Each step has hand-crafted [80, 8, 89, 3, 24, 17] and learned variants [92, 78, 93, 25, 98] and some methods operate in an end-to-end manner [78, 93, 15]. One performance bottleneck is the point classification that gets affected by the noise in the point cloud due to laser measurement noise or SfM errors. Another challenge is that the 3D point classification is heavily imbalanced: there are many more non-edge points than edge ones. These challenges are typically addressed by including point filtering steps [14], weighted losses [93, 78] and robust edge parameter estimation [19, 69]. An efficient alternative to the 3D edge point classification is to work on a 3D edge point cloud directly, i.e., a point cloud with points only on the edges of the scene.
@@
Recent works [91, 87, 44] demonstrate how to obtain such 3D edge point clouds from images: they train neural implicit fields that represent the 3D edges and supervise it with the standard rendering loss. The supervisory signals are images [87] or 2D edge maps [91, 44] generated by off-the-shelf detectors [18, 73, 62, 88]. The resulting 3D field is then sampled to get the 3D edge points on top of which the point clustering, linkage, and curve fitting are run. NEF [91] learns an edge density field that represents the probability of a 3D point to lie on a 3D edge. The point sampling then amounts to sampling the 1-level-set. Inspired by VolSDF [90], the edge density is tied to a volumetric density [54] so that NEF can be trained simply with volumetric rendering on the 2D edge maps [73]. NEF is trained with three losses to account for the specificities of 3D edge fields: a sparsity loss to encourage sparse edges, a balancing loss to account for the class imbalance between edge points and non-edge points, and a 3D view consistency loss to account for the occlusions in the supervisory signal. More specifically, NEF renderings display both the visible edges and the occluded ones whereas the 2D edge maps display only the visible ones (See Fig. 2). This consistency loss prevents NEF from wrongly penalizing the rendering when it displays the occluded edges. NEAT [87] adopts a Signed-Distance-Function (SDF) [90] to learn a 3D wireframe field from images. The 3D wireframe points lie on the 0-level set of the field. More recently, EMAP [44] learns a 3D edge field with an Unsigned-Distance-Function (UDF), which is better suited to the problem than SDF that are limited to watertight 3D structures. Compared to NEF [91], they account for the class imbalance and the view-inconsistent occlusions with importance sampling that samples rays along 2D edge and 2D non-edge pixels equally. One common limitation of these methods is their reliance on computationally heavy neural implicit representations to derive the 3D edge point cloud. Also, the point sampling is done within an $\epsilon$-bound of the level sets [91, 44] and such points do not lie accurately on the 3D edges, which require further post-processing.
Instead, we propose to learn directly the 3D edge point cloud with the explicit and extremely efficient Gaussian Splat [38] representation, which is inherently suited to represent edge point clouds: i) the 3D edge points are simply the centers of the learned Gaussians so no point sampling is required; ii) the main direction of the Gaussian defines the 3D edge point direction, which is the direction of the edge it lies on. This greatly simplifies the linkage and clustering step. iii) our method trains an order of magnitude faster than existing methods.
@@
3D Gaussians are a relevant alternative representation of 3D edge points that allows for efficient rendering-based training. After training, we revert to the traditional representation with the following mapping: a 3D edge point is the mean of the Gaussian, with the principal axis of the Gaussian as its associated direction. The principal axis of 3D Gaussian is the one with the largest scale. Once the 3D edge point cloud is obtained, the 3D edges are derived with a standard cluster-then-fit approach that first clusters the points into groups that lie on the same edge and then fits an edge to each cluster.
In this section, we first provide an overview of the original 3D Gaussian Splatting framework introduced in [38]. We then present an adaption of the splatting framework for the 3D edge reconstruction using multi-view 2D edge maps. Last, we describe how edges are fitted from the proposed explicit representation.
@@ 3.1 Preliminaries: 3D Gaussian Splatting (3DGS)
3D Gaussian Splatting [38] represents the scene using a set of 3D Gaussians. The mean $\mu\in\mathbb{R}^{3}$ of the Gaussian defines its position, and the scale and the orientation of the Gaussian are defined by its covariance matrix $\Sigma\in\mathbb{R}^{3\times 3}$. A 3D Gaussian centered at $\mu$ is defined as:
@@
Each Gaussian also has an opacity attribute $\alpha$ and a color attribute $c$ represented by spherical harmonic coefficients. All the Gaussian’s parameters are differentiable and are trained by rendering the 3D Gaussians and comparing the rendered images to the original images. The loss is a combination of the $\mathcal{L}_{1}$ loss and the Difference of Structural Similarity loss [79] (D-SSIM) $\mathcal{L}_{\text{SSIM}}$ between the two, weighted by $\lambda\in\mathbb{R}$:
@@
The rendering first projects the 3D Gaussians to 2D splats [99]. The color of a pixel in the rendering is derived by selecting the Gaussians which splats overlap with the pixel, ordering them based on their depth, and $\alpha$-blending their colors. To optimize the covariance matrix $\Sigma$ while maintaining its positive-semi definiteness, it is decomposed into a rotation matrix $R\in\mathbb{R}^{3\times 3}$ and a diagonal scaling matrix $S\in\mathbb{R}^{3\times 3}$, such that $\Sigma=RSS^{T}R^{T}$.
The representation is initialized with Gaussians centered at a sparse set of points, e.g., obtained from SfM [70]. The set of Gaussians in the scene is dynamically controlled by duplicating, splitting, and culling Gaussians based on various criteria - the important ones being 1) culling low-opacity Gaussians; and 2) duplicating/splitting Gaussians that are in a region that needs densification, or is a complex 3D surface requiring more Gaussians.
Figure 2: Occlusions for 3D edges. The red edges are the occluded edges absent from the supervisory 2D edge maps. This occlusion is inherited from the occlusion in the natural image (left) from which the 2D edge maps are inffered [73, 62]. Yet, these edges are present in the rendering of the 3D edge representation, which is the desired behavior.
@@ 3.2 Gaussian Splatting for Edge Images
The original Gaussian Splatting previously described is designed to reconstruct scenes in the wild so that novel views can be accurately and efficiently rendered. It is agnostic to the content of the images so it can train on 2D edge maps [18, 73, 62, 88] in a straightforward manner. Such an off-the-shelf training produces reasonable renderings. However, the properties of 3D edges, such as their geometry and their sparsity, and the inconsistencies between the 3D Gaussian renderings and the 2D edge maps used as supervisory signals require some adaptation of 3DGS, which we describe next.
Before, note that 3D edges are colorless by nature so we discard the spherical harmonic parameters used to model colors and view-dependent appearance, which significantly reduces the number of parameters [21] and contributes to our method’s efficiency.
Figure 3: The two geometric criteria for the learned 3D Gaussians. Given the 3D edge point cloud, the 3D edges are derived by first clustering the points that belong to the same edges and then fitting an edge to each cluster. To enable such clustering, the Gaussians associated with two 3D edge points that lie on the same edge and that are neighbors must be spatially close and have similar directions (left) (see Sec. 3.2). Gaussians without the direction constraints are illustrated on the right. The white pixels of the grid here show edge pixels while grey ones show background.
@@
3D edge sparsity and incomplete supervision.  Natural images typically depict the surfaces of the scene, which is a dense 3D geometry, so there are relatively few empty spaces. This is the opposite of 2D edge maps that depict 3D edges that are inherently sparse. As already observed in previous works [44, 91], such an extreme imbalance in the pixel distribution can hinder the training convergence. For example, since edge maps consist mostly of zero-value pixels where no 3D edge reprojects and few pixels with non-zero values, the training could converge to valid but degenerate sets of Gaussians such as an empty set of Gaussians or Gaussians with only background color.
@@
Another issue lies in the supervisory signal: the proposed method, as well as previous works [91, 44], supervise their training with 2D edge maps generated by off-the-shelf detectors [73, 62]. The edge maps are generated from natural images with occlusions that the edge maps inherit even when the occlusions do not affect the 3D edges. We illustrate this phenomenon in Fig. 2: the supervisory edge map is generated from the natural image (left) and comprises only the visible edges (black edges on the right). However, the rendering of the 3D edges learned by the trained representations do show the occluded edges (red), which is the desired behavior. Naively training the model with the standard rendering loss would penalize the model for rendering the occluded edges, i.e., the red edges.
To address these issues, NEF [91] weights the rendering loss based on the pixel’s probability to be an edge, i.e., the pixel’s intensity in the edge map. This penalizes the edge pixels more than the non-edge pixels for which the supervisory signal is less reliable. However, this reduces the supervisory signal on occluded edges and the trained Gaussians would result in incomplete reconstructions: the 3D edges seen only by few edge maps would get discarded because their opacity would be too low. EMAP [44] works around this limitation with importance sampling that samples rays uniformly between edge and non-edge pixels. We adopt a similar strategy except that we sample uniformly at random between edge and non-edge pixels instead of rays to compute the projection loss.
@@
This is derived by simply masking the $\mathcal{L}_{1}$ loss between a rendered image $\hat{I}$ and a training image $I$ as follows:
@@
where $\odot$ is Hadamard product and $\mathcal{M}$ is a 2D binary mask with values 1 at all edge pixels and at an equal number of randomly selected background pixels. Also, we discard the SSIM term in Eq. 2.
Learning oriented edge points as Gaussians.  The learned Gaussians should be such that the resulting edge points not only lie on the 3D edges but also are suitable for the subsequent edge fitting. As detailed in Sec. 3.3, we adopt a cluster-then-fit strategy that first clusters the points into groups that lie on the same edge and then fits an edge to each cluster. The clustering is cast as a simple graph traversal where the vertices are the oriented edge points: the vertex’s position is the Gaussian’s center and the vertex’s direction is the Gaussian’s principal direction. To allow such clustering to identify edges, the Gaussians should comply with two criteria illustrated in Fig. 3: i) the Gaussians belonging to the same edge should have centers that are spatially close; ii) the Gaussians belonging to the same edge should have a similar direction, i.e., their principal axes should be colinear. The second criterion is enforced by constraining the principal direction of the Gaussian with the largest axis to align with the direction of the edge it lies upon. A simple way to implement such a constraint is to encourage the principal axis of a 3D Gaussian to point towards its $k$ nearest neighbors in 3D, $k$ being a hyperparameter. Mathematically, we enforce the magnitude of the dot product between the principal axes of neighboring Gaussians to be close to 1. Given $N$ 3D Gaussians indexed between $1$ and $N$, let $\mu_{i}$ and $d_{i}$ be respectively the center and the principal direction of the $i^{th}$ Gaussian, and let $i_{1},...,i_{k}$ be the indices of the neighbors, we enforce the constraint with the following loss:
@@
We also regularize the Gaussians to have an ellipsoidal shape instead of spherical or disc-like shape so that identifying the principal direction of the Gaussian, i.e., the largest axis, is computationally more stable. For example, in the extreme case where the Gaussian is a perfect sphere, all directions would be the principal directions since they would have the same scale. The regularization encourages the largest axis of the Gaussian to be distinctive enough, which amounts to enforcing the ratio between the largest scale and the next one to be large:

### v2

@@ Abstract
With their meaningful geometry and omnipresence in the 3D world, edges are extremely useful primitives in computer vision. Methods for 3D edge reconstruction have 1) either focused on reconstructing 3D edges by triangulating tracks of 2D line segments across images or 2) more recently, learning a 3D edge distance field from multi-view images. The triangulation-based methods struggle to repeatedly detect and robustly match line segments resulting in noisy and incomplete reconstructions in many cases. Methods in the latter class rely on sampling edge points from the learnt implicit field, which is limited by the spatial resolution of the voxel grid used for sampling, resulting in imprecise points that require refinement. Further, such methods require a long training that scales poorly with the size of the scene. In this paper, we propose a method that explicitly learns 3D edge points with a 3D Gaussian Splatting representation trained from edge images. The 3D Gaussians are regularized to have their directions of largest variance along the edge they lie on, enabling clustering into separate edges. Backed by efficient training, the proposed method produces results better than or at-par with the current state-of-the-art methods, while being an order of magnitude faster. Code released at https://github.com/kunalchelani/EdgeGaussians.
Figure 1: EdgeGaussians: the proposed method learns oriented 3D edge points via Gaussian Splatting specialized for 2D edge maps. The mean and the direction of largest variance of a Gaussian define an edge point’s position and orientation. Left to right: 2D edge maps generated by off-the-shelf detectors [66, 57] are used as supervisory signals to train the 3D edge Gaussians. The trained Gaussians are clustered based on spatial proximity and orientation consistency. Parametric edges are fitted on top of these clusters.
@@ 1 Introduction
3D edges comprise of straight 3D lines and 3D curves, which we will refer to respectively as lines and curves for simplicity. Seminal works reconstruct 3D lines from images with a Structure-from-Motion (SfM) approach where line segments are detected from images, matched, and triangulated into 3D lines [74, 75, 5, 52, 6, 58]. The approach extends to 3D curves [32, 62, 60, 33]. However, the lack of repeatability in the 2D edge detection and the limited robustness in the edge matching remain performance bottlenecks for such approaches. although recent works have made impressive progress in mapping lines [56, 55, 43].
@@
3D edge extraction from 3D point clouds is free of those limitations and usually follows three steps: classify which 3D points lie on edges, cluster the points belonging to the same edge, and link them [73, 8, 84, 71]. However, the classification is usually hindered by the extreme imbalance between the edge point and the non-edge points, and noisy point clouds can lead to spurious classification. An efficient alternative is to operate directly on a 3D edge point cloud, i.e., a point cloud with points only on the 3D edges.
@@
Recent methods [79, 41, 83] have proposed to sample such an edge point cloud from neural implicit fields learned to represent 3D edges using multi-view images or 2D edge maps [16, 66] as supervision. Parametric 3D edges are then fit on these sampled edge points. However, such neural fields are computationally expensive and require long training times: recent methods[83, 79, 41] take several hours to train on a simple CAD model from the ABC [83, 38] dataset (see Sec. 4 for precise runtimes). Another limitation is the accuracy at which neural implicit fields can represent 3D edges: in theory, the 3D edge points lie on a level set of the field (0 for distance fields [41, 79] and 1 for probability fields [83]). In practice though, the field is evaluated at a finite 3D resolution and sampling points at the exact level set is not feasible. To compensate for this, the point sampling is done within an $\epsilon$-bound of the level sets [83, 41] but such points do not lie accurately on the 3D edges, requiring post-processing to correct these errors.
@@
To make the 3D edge reconstruction simpler and more efficient while preserving accuracy, we propose to learn an explicit representation of the 3D edge points. Our method directly learns an edge point cloud, bypassing the need for level-set point sampling and post-processing. It also directly learns the edge direction at each point, instead of requiring an additional step to infer it [41]. The associated edge direction to each point makes the 3D edge fitting simpler. Last, the method is several times faster (30 and 15 times as compared to  [41] and  [83], respectively), while producing results at-par with or slightly better than the current state-of-the-art. The proposed method optimizes 3D Gaussians to have their means close to the 3D edges and their direction of largest variance to be along the direction of the edge. These 3D Gaussians are then mapped to oriented points. Such a representation is geometrically meaningful and is trained in a fast and straightforward manner by adopting the training optimization defined in 3D Gaussian Splatting (3DGS) [34]. The optimization is adapted to accommodate the specificities of 3D edge learning mainly the sparsity and the occlusion of the 3D edges. The training remains simple and is supervised with off-the-shelf 2D edge maps [66, 57], as in [41, 83].
@@ 2 Related work
We first review the 3D edge reconstruction methods based on multi-view images - the seminal SfM-based methods and the more recent ones that learn implicit neural edge fields. We then discuss the methods for extracting 3D edges from point clouds.We also review 3D Gaussian Splatting methods, especially those with geometry constraints.
@@
3D Edges from multi-view images.  Traditional methods reconstruct 3D lines from images with an SfM approach that detects lines [70, 55, 28, 78, 2], matches them across images based on line descriptors [56, 39, 69, 7], and lifts them to 3D with triangulation [75, 74, 88, 5, 52, 6, 13, 87, 49, 43, 58, 61] or epipolar geometry [23, 24, 25]. The main challenges are repeatedly detecting lines even if they are occluded and matching lines robustly across images. While recent works [43, 23, 24, 25, 22] have made impressive progress in addressing these limitations, the robustness of the line detection and matching remains a performance bottleneck. These limitations also hold for similar methods that reconstruct 3D curves [32, 62, 60, 33]. Alternative approaches work around line matching by estimating 3D lines using geometric graph optimization on the detected 2D line segments [30, 59], or directly predicting the 3D wireframe in an end-to-end learned manner [90, 46] but they fall behind SfM-based method in reconstruction accuracy [43]. In this paper, we propose a method that is on-par with the state-of-the-art without relying on edge detection and matching.
@@
Recent works [83, 79, 41] avoid the limitations of SfM-based methods by learning a neural implicit field to represent the 3D edges. The supervision for training such a field are natural RGB images [79] or 2D edge maps [83, 41] generated by edge detectors [16, 66, 57, 80]. The resulting 3D field is then sampled to get the 3D edge points on top of which point clustering, linkage, and curve fitting are performed. NEF [83] learns an edge density field that represents the probability of a 3D point to lie on a 3D edge. The point sampling then amounts to sampling the 1-level-set. Inspired by VolSDF [82], the edge density is tied to a volumetric density [50] so that NEF can be trained simply with volumetric rendering on the 2D edge maps [66]. EMAP [41] improves on NEF [83] and produces state-of-the-art results by learning a 3D edge field with an Unsigned-Distance-Function (UDF). Both methods account for the class imbalance and the view-inconsistent occlusions of 2D edge maps with weighted rendering loss [83] and importance sampling of rays  [41]. NEAT [79] adopts a Signed-Distance-Function (SDF) [82] to learn a 3D wireframe field from images jointly with junction points to derive the wireframes. One common limitation of these methods is their reliance on computationally heavy neural implicit representations resulting in long training times. Also, the edge points are sampled over level sets approximated with voxel grids so, even with a fine resolution, such points do not lie accurately on the 3D edges and require further refinement. Instead, our method explicitly learns oriented points along the 3D edges in the form of 3D Gaussians. These oriented points are then clustered into individual edges which are then represented in parametric form.
3D Edge extraction from point clouds.  To extract edges from a set of 3D points, most approaches first classify which 3D points lie on edges, cluster the points that lie on the same edge, link them, and fit a parametric edge to each cluster. Each step has hand-crafted [73, 8, 81, 3, 20, 15] and learned variants [84, 71, 85, 21, 91] and some methods operate in an end-to-end manner [71, 85, 14]. Such methods generally require a dense and noise-free point cloud as input and do not directly work on SfM point clouds obtained from multi-view images.
3D Gaussian Splatting.  The seminal work of [34] introduces the representation of scenes as 3D Gaussians which can be rendered efficiently. The parameters of Gaussians are learned using a loss function that compares the rendered and ground-truth images from multiple views. Similar to our approach, certain works introduced geometric regularization in 2D or 3D to model human hair strands using 3D Gaussians [86, 45]. However, they have been applied and tested specifically for the task of hair modeling.
@@
Figure 2: Occlusions for 3D edges. The red edges are the edges occluded by surface and are absent from the supervisory 2D edge maps. Yet, these edges are present in the rendering of the 3D edge representation, which is the desired behavior.
Given a set of edge images, our method directly derives oriented points along 3D edges. These points are then clustered into individual edges, on which parametric fitting is performed. A relevant learnable representation of an oriented 3D edge point is a 3D Gaussian; with the mean of the Gaussian being the point’s position and its principal direction being the point’s orientation. Such a representation is efficiently trained using the 3DGS [34] optimization with edge images as input. In this section, we first review the original 3DGS [34] framework. We then discuss how 3DGS is adapted for 3D edge reconstruction from multi-view 2D edge maps, and lastly, we describe the parametric edge fitting from the optimized 3D Gaussians.
@@ 3.1 Preliminaries: 3D Gaussian Splatting (3DGS)
3D Gaussian Splatting [34] represents the scene using a set of 3D Gaussians. A 3D Gaussian centered at $\mu$, with covariance $\Sigma$ is defined as:
@@
Each Gaussian also has an opacity attribute $\alpha$ and a color attribute $c$ represented by spherical harmonic coefficients. All the Gaussian’s parameters are differentiable and are trained by rendering the 3D Gaussians and comparing the rendered images to the original images. The loss is the sum of the $\mathcal{L}_{1}$ loss and the Difference of Structural Similarity loss [72] (D-SSIM) $\mathcal{L}_{\text{SSIM}}$ between the two, weighted by $\lambda\in\mathbb{R}$:
@@
The rendering first projects the 3D Gaussians to 2D splats [92], and the color of a pixel is then derived by $\alpha$-blending the colors of the Gaussians projecting on the pixel. To optimize the covariance matrix $\Sigma$ while maintaining its positive-semi definiteness, it is decomposed into a rotation matrix $R\in\mathbb{R}^{3\times 3}$ and a diagonal scaling matrix $S\in\mathbb{R}^{3\times 3}$, such that $\Sigma=RSS^{T}R^{T}$.
The representation is initialized with Gaussians centered at a sparse set of points, e.g., obtained from SfM [63]. The set of Gaussians in the scene is dynamically controlled by duplicating, splitting, and culling Gaussians based on various criteria - the important ones being 1) culling low-opacity Gaussians; and 2) duplicating/splitting Gaussians that are in a region that needs densification, or a complex 3D surface requiring more Gaussians.
@@ 3.2 Gaussian Splatting for Edge Images
3DGS [34] is designed for fast and accurate novel view synthesis while our focus here is more on positioning of the Gaussians along 3D edges. Additionally, the supervision in our method comes 2D edge maps instead of natural RGB images. These differences in the final objective and the data modality motivate certain changes in the training paradigm. We first detail the issues that arise when using edge images for supervision and propose a solution to address these issues. Then we present a geometric regularization to steer the training objective from purely view synthesis, towards one focused on the geometry of 3D Gaussians useful for edge modeling.
Figure 3: Geometric regularization of 3D Gaussians. Enforcing elliptical Gaussians with the principal direction aligned with the neighbors would result in 2D projections as shown on the left, while un-regularized ones might look like those on the right. The ones on the left are geometrically meaningful and result in easier clustering into separate edges.
@@
Supervision from edge maps.  Natural images typically depict the surfaces of the scene, which is a dense 3D geometry. On the other hand, 2D edge maps depict edges that are inherently sparse. As already observed in previous works [41, 83], such an extreme imbalance in the pixel distribution can hinder the training stability. For example, since edge maps consist mostly of zero-value pixels where no 3D edge reprojects and few pixels with non-zero values, the training could converge to valid but degenerate configurations such as an empty set or a set Gaussians with only background color.
@@
Additionally, 2D edge maps [16, 66, 57, 80] exhibit occlusions due to the scene’s surfaces that do not occur when rendering 3D edges. This makes the supervision signal incomplete as illustrated in Fig. 2: the supervisory edge map is generated from the natural image (left) and comprises only the visible edges (black edges on the right). However, the rendering of the 3D edges learned by the trained representations do show the occluded edges (red), which is the desired behavior. Naively training the model with the standard rendering loss would penalize the model for rendering the occluded edges, i.e., the red edges.
To address this issue, previous works [83, 41] weight the rendering loss higher for edge pixels as compared to non-edge pixels [83], or compute the loss over pixels sampled uniformly across the edge and non-edge pixels [41]. In our experiments, we observe that the former is more efficient while the latter leads to slightly better results, which we adopt and detail next.
@@
The sampling is derived by masking the $\mathcal{L}_{1}$ loss between a rendered image $\hat{I}$ and a training image $I$ as follows:
@@
where $\odot$ is the Hadamard product operation and $\mathcal{M}$ is a 2D binary mask with values 1 at all edge pixels and at an equal number of randomly selected background pixels. Also, we discard the SSIM term in Eq. 2.
Learning oriented edge points as Gaussians.  The 3D Gaussians optimized only with the rendering loss can be oriented and scaled without correlation to the 3D edge geometry. This impedes the subsequent edge fitting so we introduce geometric regularization illustrated in Fig. 3. As later detailed in Sec. 3.3, we adopt a cluster-then-fit strategy that first clusters the points into groups that lie on the same edge and then fits an edge to each cluster.
Clustering edges only based on spatial proximity raises ambiguities when clustering points close to two or more edges - points close to junctions. A standard way to address such ambiguity is to cluster points that are not only spatially close but also have the same orientation [53]. To enable this, the learned edge points must have the same orientation as the edge they lie on. This is equivalent to enforcing the principal direction of each 3D Gaussian, i.e., the direction of its largest variance, to align with the edge’s direction. In practice, we implement such a constraint by encouraging the principal direction of a 3D Gaussian to point towards its $k$ nearest neighbors in 3D, $k$ being a hyperparameter.
Given $N$ 3D Gaussians indexed between $1$ and $N$, let $\mu_{i}$ and $d_{i}$ be respectively the center and the principal direction of the $i^{th}$ Gaussian, and let $i_{1},...,i_{k}$ be the indices of the neighbors, we enforce the constraint with the following loss:
@@
Note that this approach works in general when the $k$ nearest neighbors lie on the same edge and are optimized to approximately correct positions. However, it can break down in the presence of fine structures as shown in the failure cases.
We also regularize the Gaussians to have an ellipsoidal shape instead of a spherical or disc-like shape so that identifying the principal direction of the Gaussian, i.e., the largest axis, is computationally more stable. This amounts to enforcing the ratio between the largest scale and the next one to be large: